OpenStudy (anonymous):

Which of the following equations has the solution set {0}? -14m- 7m+ 1 = -7m+ 1 -14m+ 7m+ 1 = -7m+ 2 14m- 7m- 1 = -7m+ 1

OpenStudy (cwrw238):

just plug m = 0 into each equation and see if both sides of '=' is the same

OpenStudy (anonymous):

It is the first one: after simplifying it it is: -14m = 0 => m = 0; 2nd one is: 1-7m = 2 - 7m 3th give you result of m = 1/7